### v1

# Predicting the adhesion and delamination strength of carbon films on metals by high-throughput ab initio calculations

Abstract
Diamond and diamond-like carbon (DLC) coatings are widely employed for their exceptional mechanical, thermal and chemical properties, but their industrial application is often limited by weak adhesion to metallic substrates. In this work, we employ a high-throughput ab initio approach to systematically investigate the adhesion of diamond/metal interfaces, combining a set of technologically relevant metals (Al, Ag, Au, Cr, Cu, Fe, Ir, Mg, Mo, Pt, Rh, Ti, V, W, Zn) with the C(111), C(111)-21 (Pandey reconstructed), C(110), C(100), that are most common in diamond and are representative of different types of bonds present in DLC. Thanks to our automated and accurate computational protocol for interface construction and characterization, databases are populated and relevant trends are identified on the effect of surface graphitization, ability to form carbides and metal reactivity on carbon film adhesion and delamination strength. Beyond capturing trends, our workflow yields predictive insights. Indeed, we found that adhesion energy scales with the geometric mean of the constituent surface energies, providing a simple descriptor for rapid screening; while comparing the work of separation with the metal’s cohesive energy anticipates the fracture location under tensile loading. A novel method based on analysis is introduced to identify when contact with a metal drives rehybridization of surface carbon from sp2 to sp3, the structural signature of improved resistance to delamination. These structural changes are mirrored by an electronic rearrangement at the interface, quantified by a charge-accumulation descriptor that strongly correlates with adhesion.
I INTRODUCTION
Diamond displays unmatchable properties: extreme hardness, chemical inertness and biocompatibility, exceptionally high thermal conductivity, low dielectric constant, high thermal stability and excellent resistance to radiation. Due to these qualities carbon-based films, made of amorphous diamond like carbon (DLC) and poly- or nano-crystalline diamond, have found application within the development of sensors, electrodes, high power and high temperature devices and substrates for biological activity able to comply also with harsh environments [1, 2, 3, 4]. Besides these qualities, diamond and DLC films coatings provide some of the lowest known wear and friction coefficients [5, 6] and are therefore applied in industry to minimize friction and wear in engine parts, prevent stiction in micro- and nano-electromechanical systems (MEMS and NEMS) and enhance durability and performance of industrial cutting tools and biological implants [7, 8, 9, 10].
Being able to control adhesion of a DLC film by tuning the parameters that govern it, is pivotal also within the DLC film itself. Indeed, the weak adhesion of carbon coatings on metallic substrates has long been one of the main obstacles to their widespread industrial application, as it can eventually lead to spallation or delamination, making the coating entirely ineffective [11, 12, 13]. Conversely, a weak adhesion with the countersurface is crucial to achieve low friction and wear [14]. For these reasons to fully exploit the potential of DLCs, it is essential to carefully design and optimize the combination of substrate, coating and countersurface and, ideally, to tune those properties that govern adhesion while growing the film itself so that it proves ”sticky” on one side and unreactive on the other.
One of the most established strategies to improve adhesion with the substrate is to modify the substrate itself through the introduction of a metallic interlayer, which helps to relieve residual stress within the coating and provides a strong chemical bond between carbon and metal substrate atoms. This is particularly important for steel or iron substrates, where carbon diffusion during chemical vapor deposition (CVD) can lead to the formation of graphite-like layers that weaken the interface bonding [15, 16, 17]. Both theoretical [18, 19] and experimental [6, 20, 21, 22] studies have shown that carbide-forming metals, such as Ti, Cr and W, are highly effective as interlayers, ensuring the adhesion required for DLC to maintain its anti-wear functionality.
At the same time, the nature of the carbon coating itself also strongly influences its tribological properties. A higher fraction of sp2 hybridized carbon atoms leads to a more graphite-like character, producing a coating with low friction coefficient, whereas a larger amount of sp3 carbons leads to a diamond-like structure with superior hardness and elastic modulus [23, 24]. Hence, it is clear that the sp3/sp2 ratio is a crucial factor governing adhesion and wear behavior of amorphous carbon coatings and should not be overlooked [15, 25, 26, 27].
Given the numerous factors that influence the performance of carbon-based coatings, i.e., the type of substrate and interlayer material, the properties of the DLC film and those of the countersurface, experimental testing of all possible combinations just to find the best one for a specific application is extremely time and resources consuming. This is where accurate and systematic ab initio calculations could play a role to help speed up the process through microscopic understanding and accurate datasets of theoretical predictions. Indeed, such a systematic study enables the rapid identification of optimal candidates for specific applications to be further tested experimentally, while also revealing quantitative relationships between adhesion and other fundamental physical quantities, as will be discussed in the following sections. Nevertheless, theoretical studies on diamond/metal interfaces remain limited, both in terms of the kind of metals and surface orientations investigated and in the diversity of the employed computational setup and system modeling [18, 19, 28, 29, 30, 31]. This makes systematic comparisons difficult and leaves the factors governing adhesion and tribological performance poorly understood.
In this work, we address this gap using an ab initio high-throughput approach, enabling a systematic evaluation and comparison of adhesion energies across technologically relevant diamond/metal interfaces. Furthermore, we propose a rigorous computational protocol for constructing the interface supercell. In contrast to the common practice of simply adapting the lattice of one component to the other, typically by straining the metallic substrate to match the diamond lattice often resulting in unphysical deformations, our approach employs the algorithm developed by Zur, which identifies the smallest common supercell between two crystalline surfaces while controlling lattice mismatch and angular distortion below user-defined limits [32]. A detailed description of the algorithm and its implementation is provided in Section II.
On the metal side, this study considers the most stable terminations of Al, Ag, Au, Cr, Cu, Fe, Ir, Mg, Mo, Pt, Rh, Ti, V, W and Zn; while four different low-index diamond surfaces are chosen to represent the different types of bonds in DLC coatings: the non-reconstructed C(111)(11) surface with dangling bond (DB) termination, its most stable reconstruction C(111)(21) called Pandey reconstruction, the C(110)(11) surface and the so-called Dimer reconstructed C(100)(21) termination. From now on, we will address them as C(111), C(111)Pd, C(110) and C(100), respectively. C(111) is the only one exhibiting a sp3 bond structure at the surface; while C(111) Pd, with its zig-zag surface chains and delocalized electrons, is representative of the sp2 aromatic or graphitic layer commonly observed on the diamond/DLC surface during rubbing. C(110) also forms zig-zag -bonded chains, but its stabilization and inertness is less pronounced than that of C(111)Pd; whereas C(100) stabilizes through surface dimers, resulting in a distorted sp2-like geometry [33, 34, 35].
The paper is organized as follows: in Section II, we describe the computational protocol implemented in our high-throughput software for the automated construction and characterization of diamond/metal interfaces. Section III presents and discusses the results, focusing first on the trends in adhesion energies across the various diamond terminations and metal substrates and on how these relate to the intrinsic properties of the constituents, such as their surface energy. Then, the fracture behavior of the interfaces is examined to assess their mechanical stability under tensile load. Finally, electronic and structural responses upon interface formation, such as charge accumulation and modification of diamond surface reconstruction, are addressed, providing a microscopic understanding of the adhesion mechanisms.
II METHODS
The adhesion of diamond/metal interfaces has been systematically investigated through automatized simulations using TribChem [36], an advanced software for the high-throughput atomistic study of solid-solid interfaces and of their tribological properties, developed by our group. TribChem is written entirely in Python and performs Density Functional Theory (DFT) simulations using the Vienna Ab initio Simulation Package (VASP) [37, 38, 39]. The core of the code is the FireWorks workflow manager [40], which, in addition to executing workflows, allows both the creation of custom workflows and the use of the built-in ones available in the Atomate package [41]. In Tribchem, most workflows are customized, but Atomate is still employed, especially to manage VASP runs and recover errors through its built-in integration with the Custodian package [42]. Other essential libraries are Pymatgen [43] and MPInterfaces [44], used for handling input as well as pre- and post-processing operations.
A key feature of TribChem is its systematic storing of relevant physical results. In general, the database is a fundamental component in high-throughput workflows; however, in TribChem, beyond the internal FireWorks database used for tasks management, a new high-level one is created to store physical quantities returned by the workflow in an organized manner. In particular, we use MongoDB, a NoSQL database, which facilitates future data analysis and ensures that data are not recomputed.
The software has been well tested for homogeneous and heterogeneous metal-metal interfaces [45, 46], and it has been recently extended to handle the more complex structure of semiconducting materials.
Tribchem allows for the automatic calculation of several key parameters, including bulk modulus, surface energy, adhesion energy and charge transfer during interface formation, in order to analyze structural and tribological characteristics of these systems.
The workflow starts by reading the user-defined input parameters, comprising the name of the workflow to be executed, the unique identifiers of the basic bulk structures and the Miller indices that identify the slab orientation for each material. Additional parameters, such as the DFT exchange and correlation functional, whether Van der Waals corrections should be used, the maximum allowed cross-section area of the interface supercell and many others, can be supplied, otherwise default values are retrieved from a central JSON file. In this work, the Perdew-Burke-Ernzerhof generalized-gradient approximation (PBE-GGA) [47] of the exchange-correlation functional has been employed and all calculations are spin polarized.
A schematic representation of the computational workflow is shown in Fig. 1. The first step consists in retrieving the input structures for the selected materials from the online Materials Project database [48] and storing them in our database. Both bulk structures are then optimized in parallel, by converging their lattice parameters via a fit with the Birch-Murnaghan equation of state [49] and optimizing the kinetic energy cutoff and k-point density mesh required by DFT. The higher values for energy cutoff and k-point density are chosen after all convergence loops have finished, so that the interfaces calculations will be both efficient and accurate. Subsequently, slabs are obtained by cutting the bulk along the selected directions and their thickness is converged with respect to the surface energy. The physical quantities obtained from the bulk and surface workflows for the selected metals and diamond terminations are reported in Tab.1; while optimized computational parameters are shown in Tab.S1. The diamond surfaces characterized by complex reconstructions, namely C(111)Pd and C(100), were constructed outside the automated workflow, but still using the bulk parameters optimized by TribChem.
| Material | (Å) | K (GPa) | () | (J/m2) |
| W | 2.759 | 302 | (110) | 3.28 |
| Cr | 2.866 | 179 | (110) | 3.19 |
| Mo | 2.736 | 258 | (110) | 2.83 |
| Fe | 2.452 | 179 | (110) | 2.51 |
| V | 2.597 | 182 | (110) | 2.40 |
| Ir | 2.738 | 348 | (111) | 2.31 |
| Rh | 2.708 | 255 | (111) | 2.03 |
| Ti | 2.935 | 112 | (001) | 2.02 |
| Pt | 2.806 | 247 | (111) | 1.49 |
| Cu | 2.569 | 140 | (111) | 1.29 |
| Al | 2.860 | 77 | (111) | 0.84 |
| Ag | 2.934 | 90 | (111) | 0.62 |
| Au | 2.940 | 138 | (111) | 0.71 |
| Mg | 3.208 | 36 | (001) | 0.54 |
| Zn | 2.625 | 73 | (001) | 0.32 |
| C | 2.526 | 430 | (111)Pd | 3.43 |
| (100) | 4.66 | |||
| (110) | 5.15 | |||
| (111) | 5.04 |
As already anticipated in Section I, the matching of the two slabs to construct the heterointerface is performed following the Zur algorithm [32], implemented in the MPInterfaces library, which constructs the smallest supercell possible while satisfying specific criteria on lattice dimensions and angles strains. Further details on the algorithm are thoroughly discussed in Ref.[46]. The magnitude of the applied strain is scaled according to the inverse of the bulk modulus of each material, meaning that materials with higher compressibility undergo greater deformation during interface construction. However, in this case, since diamond is significantly stiffer than the metals considered, the strain was applied exclusively to the metallic component. Since lattices with different symmetries must be accommodated within a single supercell, the strain on the metal is anisotropic in some cases. Fig.2 reports the maximum percentage strain, while average strain values are provided in Fig.S1. For all metals, the lattice strain is below 3, which significantly reduces the impact of mechanical stress on the interface electronic properties. Details on the transformations applied to the slab cells to construct the interface supercell, as automatically computed by TribChem, can be found in the Supporting Information (SI).
Finally, the tribological properties of the interface such as the Potential Energy Surface (PES), adhesion energy and charge transfer at the interface are calculated. The PES maps the interaction energy as a function of the relative lateral position of the two slabs. It is therefore obtained by relaxing the slabs in the direction normal to the interface plane for all the relative lateral shifts obtained by matching the two surfaces’ high symmetry points. Adhesion energy is then computed as:
namely by subtracting the total energies of the relaxed single slabs, and , from the total energy of the interface in the lateral position which corresponds to the PES minimum , divided by the supercell area . Electronic charge accumulation upon interface formation () is calculated as:
where is the interface distance and is the planar average of the interface charge density difference ; being the volumetric charge density of the interface, while and those of the single surfaces.
III RESULTS AND DISCUSSION
Considering a subset of technologically relevant transition metals and including Al and Mg simple metals due to their importance in alloys, the most stable facets of Ag, Al, Au, Cr, Cu, Fe, Ir, Mg, Mo, Pt, Rh, Ti, V, W and Zn have been mated to four low-index diamond surfaces: C(111), C(111)Pd, C(110) and C(100). A total of 60 diamond/metal interfaces were optimized and their adhesion were calculated. This requires to sample the whole PES and identify the energy minimum, meaning that a total of 950 DFT calculations were performed. The atomic models of interfaces formed by the four diamond terminations in contact with representative metallic substrate are shown in Fig.4.
III.1 Adhesion energy
Adhesion energies, as calculated from Eq.1, are reported in Fig.3. In general, constituents with higher surface energies form an interface characterized by higher adhesion (bottom left corner of Fig.3), whereas diamond terminations adhere less to noble metals, such as Cu, Ag, Au, which possess completely filled d-shells.
When examining Fig.3 row by row, i.e. by varying the metal substrate while keeping the diamond termination fixed, it is possible to identify some outliers. In particular, Ti shows strong adhesion to all diamond terminations, despite its relatively low reactivity, quantified by its surface energy. This behaviour can be attributed to its pronounced carbide-forming ability compared to other metals (carbide formation enthalpy of TiC per Ti atom =–185.2 kJ/mol [18] and TiC formation energy of -0.808 eV/atom [48]). Conversely, when examining the values in Fig.3 column by column, i.e. at fixed metal substrate, diamond terminations with higher surface energy exhibit stronger adhesion across all metals. An interesting comparison can be made between C(111), which exposes surface sp3 carbons, and C(111)Pd, characterized by chains of sp2 hybridized carbon atoms. While previous work reported adhesion reduction at C/Cu due to surface graphitization [50], here we see that this behavior is general and that surface graphitization drastically reduces adhesion across all metals. This suggests that while growing the DLC film, for an effective coating that maximizes the adhesion to the substrate while reducing adhesion and friction with any countersurface, a gradient of the sp2/sp3 fraction should be built with larger sp2/sp3 ratio.
These results validate our computational protocol, as it correctly predicts that carbide-forming metals such as Cr, Ti, W and V exhibit the strongest adhesion to diamond coatings, consistent with their widespread use as interlayers between DLC and metallic substrates.
III.2 Adhesion energies from intrinsic constituent properties
To reduce the computational cost of simulating the full interface, it would be useful to establish an empirical relationship linking the adhesion energy to the intrinsic properties of the individual components. For hetergoeneous metallic interfaces, it was demonstrated that adhesion energy can be approximated by the geometric mean of the surface energies of the single constituents: [46]. This expression originates from simplified physical models, where the adhesion energy is defined as ; and being the surface energies of the two constituents and the interfacial energy, i.e., the cost of creating a unit area of interface starting from the corresponding bulk of the two materials. As reported in Ref.[51], the latter can be expressed as: , thereby obtaining the relation given above.
The effectiveness of this relationship is further confirmed for diamond/metal interfaces, as illustrated in Fig.5, where is plotted against at fixed diamond termination to account for the different nature of surface bonds. and RMSE values along with the linear fit parameters are reported in Tab.2. In all cases, the relatively high values, lying in the range 70-79, indicate that the overall trend predicted by the model is robust; while deviations in slope and intercept from the ideal values of 1 and 0, respectively, highlight that predicting adhesion energy is a complex task, governed by the interplay of multiple factors, such as chemical and structural variations at the interface.
| Slope | Intercept (J/m2) | RMSE | ||
| C(111)Pd | 0.53 | -1.30 | 79 | 40 |
| C(100) | 0.50 | -0.07 | 70 | 57 |
| C(110) | 0.68 | -0.65 | 77 | 67 |
| C(111) | 0.55 | 0.43 | 72 | 62 |
III.3 Fracture location
In the context of wear, adhesion energy of an interface plays a key role in determining coating durability. Whether fracture under tensile load initiates within the bulk metal or at the metal/diamond interface directly affects the resistance of the coating to delamination and, consequently, to wear. A practical way to assess this is by comparing the work of separation, , of the diamond/metal interface with the interlayer cohesive energy of the corresponding bulk metal, a property that can be quantified by the work of separation of the homogeneous metallic interface. If the latter is lower than the former, fracture is expected to occur inside the metal; otherwise, the interface itself becomes the weakest link.
Fig.6 shows the difference between the work of separation of the semiconductor/metal interface and interlayer cohesive energy of the metal. When C(111)Pd is involved, the fracture always occurs at the interface, reflecting the fact that surface graphitization of diamond leads to extremely weak bonds with the metal and, consequently, to very low adhesion, increasing the risk of delamination. For metals with low surface energy (/2), e.g., Mg, Zn, Al, fracture may occur within the metal itself, especially when in contact with the highly reactive C(111) surface. On the other hand, metals with partially filled d-shells (W, Cr, Mo, Fe) exhibit extremely high cohesion energies, which prevent failure from taking place within the metal phase.
III.4 Structural and electronic modifications of diamond reconstructions upon adhesion to metals
To gain insight into the structural modifications of diamond reconstructions upon interface formation, we analyzed the radial distribution function of the carbon atoms. The describes the weighted probability of finding a pair of atoms at a given distance , providing direct information on the local bonding environment [52, 53, 54]. This method is widely exploited in molecular dynamics simulations to investigate short- and medium-range order in both crystalline and amorphous systems, as it offers a powerful way to capture structural rearrangements [55, 56]. In the present case, monitoring changes in the permits to detect how the reconstruction of the diamond surface evolves upon contact with the metallic counterpart, a key aspect since geometric rearrangements are known to influence the surface electronic structure [57] and therefore its adhesion strength.
The of the standalone C(111)Pd surface is shown in Fig.7(a). Peaks corresponding to the interatomic distances within the surface atoms are indicated by vertical dashed lines and illustrated in the adjacent schematic. Remaining peaks correspond to the bulk diamond structure. Here the focus is on the peak at 1.44 Å, which corresponds to the distances between atoms in the surface zig-zag chain. The area under a peak is directly related to the average number of atoms in the interval [, ] surrounding a given atom, i.e., the number of bonds per atom contributing to that peak, namely [58]:
where is the average atomic density (total number of atoms over cell volume). For C(111)Pd, the integral of the peak at 1.44 Å therefore provides a measure of the number of surface sp2 bonds per atom.
To assess the impact of metal contact on the surface graphitization, Fig.7(c) shows a qualitative comparison of the of the pristine C(111)Pd surface (plotted in the background with faded lines) with that of the same surface but within the interface geometries (black line). Different types of structural modifications are observed depending on the metal: in some cases, the peak intensity decreases without significant shift or distortion (Mo, W, Ti, Fe), while in others, dimerization (Cr, Al, Cu, Pt) or a shift in (Rh, V, Zn) is evident, reflecting a more pronounced perturbation of the surface reconstruction. Quantitatively, integrating the peak provides a measure of how many sp2 bonds remain after interaction with each metal. For example, the weak interaction with Au leaves the peak essentially unaltered, showing only minor broadening and indicating that all sp2 bonds are preserved. In contrast, Cr, which exhibits strong adhesion with C(111)Pd, induces partial dimerization of the chain and reduces the number of sp2 bonds by roughly 50. The adhesion energy of all C(111)Pd/metal interfaces as a function of the fraction of sp2 bonds preserved relative to the pristine surface is plotted in Fig.7(b). Less reactive metals, i.e., those with completely filled outer electron shells (Ag, Au, Zn, Mg), preserve the original number of sp2 bonds of the C(111)Pd surface, explaining their weak adhesion. On the contrary, more reactive metals are effectively able to perturb the surface sp2 network, leading to stronger adhesion. This behavior is highly relevant for coating applications, where metallic interlayers are commonly introduced to improve the bonding between DLC films and their substrates. In DLC systems with a high density of surface sp2 bonds, the interaction with the metal counteracts graphitization, responsible for the weak adhesion to the substrate, by promoting the rehybridization of surface carbon atoms.
The analysis was also performed for other diamond terminations exhibiting surface sp2 bonds, namely C(100) and C(110) (see Fig.S2 and S3). In general, interaction with the metallic surfaces does not lead to major perturbations of the peak. For C(100), the main effect is the reduction in peak intensity, accompanied by minor shifts in the case of Rh, V and Ir. Similarly, C(110) generally exhibits only a decrease in peak intensity, with more noticeable effects observed for Zn, showing a significant peak shift, and Mg, where dimerization of the surface chains occurs. Quantitatively, both terminations experience an overall reduction of around 50 of the original surface sp2 bonds, with the exception of C(110)/Au and C(110)/Ag, where 54 and 59 of the original bonds are preserved, respectively.
The rehybridization of the C(111)Pd surface chains is clearly reflected in the redistribution of electronic charge upon contact with the metal. As examples, we consider the interfaces with Cr and Ag, representing two extreme cases: the Cr interface exhibits dimerization of the surface chains and strong adhesion ( J/m2), while the Ag interface shows only a slight broadening of the peak, corresponding to weak adhesion ( J/m2). The volumetric charge difference, , shown in Fig.8(b) and (d), reveals that the charge rearrangement at the Cr interface is significantly more complex than at Ag. This is further quantified by the line profiles of along the direction, reported in Fig.8(a) and (c). In both cases the metal just accumulates charge in the interfacial region, while the surface carbon atoms reorganize their electronic charge density. Importantly, for the interface with Cr, this reorganization is far more pronounced: charge is removed at the surface carbon atom atomic plane where the C-C double bonds of the Pandey chain laid, moving towards the interface region, in line with the observed dimerization of the surface chains.
III.5 Charge density accumulation at the interface
The structural modifications of the diamond reconstructions directly affect the electronic charge redistribution at the interface. Previous studies have shown that the redistribution of electronic charge upon contact, quantified by the descriptor (Eq.2), correlates well with adhesion across a wide variety of interfaces, from weakly interacting Van der Waals systems (e.g., graphene) to covalently bonded materials such as diamond [59], as well as heterogeneous metallic junctions [46].
In the present case, the semiconductor–metal interaction introduces additional complexity, as surface carbon chains can undergo rehybridization and other structural rearrangements, modifying the local electronic environment. Despite this, a strong correlation is still observed between adhesion energy and , with a value of 71 (Fig.9), confirming that remains a reliable predictor even for this more intricate class of interfaces. Deviations from the linear trend arise from the interplay of chemical specificity, structural rearrangements and local orbital interactions with the metal.
IV CONCLUSIONS
In this work, a high-throughput ab initio investigation of diamond/metal interfaces has been carried out, systematically analyzing the adhesion of 60 combinations involving a broad range of metallic substrates and four low-index diamond terminations (C(111), C(111)Pd, C(110), C(100)), representative of the different types of bonds in DLC coatings. The study employs an automated workflow implemented within the TribChem software, which ensures a consistent and reproducible protocol for the construction and characterization of solid interfaces. In particular, the interface supercells were built using the algorithm developed by Zur, which identifies the smallest common supercell between two crystalline surfaces while minimizing both lattice mismatch and angular distortion. All interfaces were constructed with a metallic strain below 3, ensuring that the calculated adhesion energies are not affected by unphysical distortions.
The results show that adhesion is primarily governed by the intrinsic reactivity of the metallic component and by the type of carbon hybridization present at the surface of the diamond. Carbide-forming metals such as Cr, Ti, W and V exhibit the strongest adhesion, in agreement with their experimentally recognized role as interlayer materials that enhance DLC adhesion to metallic substrates. In contrast, noble and weakly reactive metals (Ag, Au, Cu, Zn, Mg) display poor adhesion due to their filled valence shells and low surface energies. Among the diamond terminations, the C(111) surface with sp3 bonding shows the highest adhesion, whereas surface graphitization, as observed in C(111)Pd, considerably weakens interfacial bonding.
A strong correlation between adhesion and the geometric mean of the constituent surface energies at fixed diamond termination, , confirms that adhesion strength can be rationalized from intrinsic surface properties. This provides an efficient descriptor for fast materials screening, without the need to simulate computationally expensive interface models.
By comparing the work of separation () of the diamond/metal interface with the cohesive energy of the corresponding bulk metal, it is possible to predict the likely fracture location under tensile load. Interfaces involving C(111)Pd consistently exhibit fracture at the interface, reflecting the weak bonding due to sp2 terminations. For metals with low cohesive energies, such as Mg, Zn and Al, fracture can occur within the metal bulk when in contact with highly reactive diamond surfaces, indicating that the interface is stronger than the metallic cohesion. Conversely, transition metals with partially filled d-shells (e.g., W, Cr, Mo, Fe) exhibit high bulk cohesion, preventing failure from taking place within the metal phase.
Analysis of the radial distribution function provides insights on the structural response of diamond surface reconstructions upon contact with metals. In C(111)Pd, the peak at approximately 1.44 Å, corresponding to the zig-zag surface chains, is largely preserved in the presence of weakly interacting metals such as Au and Ag, indicating minimal perturbation of the sp2 network. In contrast, Cr, Al, Cu and Pt induce significant dimerization and partial rehybridization of the surface sp2 bonds. In general, reactive metals like W, Cr, Mo, Fe, V and Ti are able to reduce the number of surface sp2 bonds by up to 50. Hence, it can be concluded that strong perturbation and reduction in the number of surface sp2 bonds in diamond coatings directly enhance adhesion, explaining why these metals are commonly used as interlayers to improve bonding between DLC films and metallic substrates.
The structural modifications are mirrored in an electronic rearrangement at the interface. Volumetric charge density differences show that metals inducing stronger structural perturbations promote more pronounced charge rearrangement, while weaker interacting metals cause only minor redistribution. Quantitatively, the charge accumulation descriptor, , remains strongly correlated with adhesion, confirming that electronic charge redistribution upon interface formation is a good predictor of adhesion strength, even in the presence of semiconductor-metal bonding.
Acknowledgements.
These results are part of the ”Advancing Solid Interface and Lubricants by First Principles Material Design (SLIDE)” project that has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant agreement No. 865633). The authors acknowledge funding by the European Union - NextGenerationEU (National Sustainable Mobility Center CN00000023, Italian Ministry of University and Research Decree n. 1033 - 17/06/2022, Spoke 11 - Innovative Materials Lightweighting) and by the European Union - NextGenerationEU (Spoke 6 - Mulstiscale Modelling Engineering Applications). The opinions expressed are those of the authors only and should not be considered as representative of the European Union or the European Commission’s official position. Neither the European Union nor the European Commission can be held responsible for them. Computational resources were provided by CINECA under the ISCRA initiative.References
- Maguire et al. [2005] P. Maguire, J. McLaughlin, T. Okpalugo, P. Lemoine, P. Papakonstantinou, E. McAdams, M. Needham, A. Ogwu, M. Ball, and G. Abbas, Mechanical stability, corrosion performance and bioresponse of amorphous diamond-like carbon for medical stents and guidewires, Diamond and Related Materials 14, 1277 (2005).
- Dearnaley and Arps [2005] G. Dearnaley and J. H. Arps, Biomedical applications of diamond-like carbon (dlc) coatings: A review, Surface and Coatings Technology 200, 2518 (2005).
- Hauert [2003] R. Hauert, A review of modified dlc coatings for biological applications, Diamond and Related Materials 12, 583 (2003).
- Wang et al. [2020] T. Wang, L. Huang, Y. Liu, X. Li, C. Liu, S. Handschuh-Wang, Y. Xu, Y. Zhao, and Y. Tang, Robust biomimetic hierarchical diamond architecture with a self-cleaning, antibacterial, and antibiofouling surface, ACS Applied Materials & Interfaces 12, 24432 (2020).
- Robertson [2002] J. Robertson, Diamond-like amorphous carbon, Materials Science and Engineering: R: Reports 37, 129 (2002).
- Lepicka et al. [2017] M. Lepicka, M. Gradzka-Dahlke, D. Pieniak, K. Pasierbiewicz, and A. Niewczas, Effect of mechanical properties of substrate and coating on wear performance of tin- or dlc-coated 316lvm stainless steel, Wear 382-383, 62 (2017).
- Scharf and Prasad [2012] T. W. Scharf and S. V. Prasad, Solid lubricants: a review, Journal of Materials Science 48, 511 (2012).
- Auciello et al. [2004] O. Auciello, J. Birrell, J. A. Carlisle, J. E. Gerbi, X. Xiao, B. Peng, and H. D. Espinosa, Materials science and fabrication processes for a new mems technology based on ultrananocrystalline diamond thin films, Journal of Physics: Condensed Matter 16, R539 (2004).
- Love et al. [2013] C. Love, R. Cook, T. Harvey, P. Dearnley, and R. Wood, Diamond like carbon coatings for potential application in biological implants—a review, Tribology International 63, 141 (2013).
- Azzi et al. [2009] M. Azzi, M. Paquette, J. Szpunar, J. Klemberg-Sapieha, and L. Martinu, Tribocorrosion behaviour of dlc-coated 316l stainless steel, Wear 267, 860 (2009).
- Wang et al. [2002] R. Wang, C. Mercer, A. Evans, C. Cooper, and H. Yoon, Delamination and spalling of diamond-like-carbon tribological surfaces, Diamond and Related Materials 11, 1797 (2002).
- Lawes et al. [2007] S. D. A. Lawes, M. E. Fitzpatrick, and S. V. Hainsworth, Evaluation of the tribological properties of dlc for engine applications, Journal of Physics D: Applied Physics 40, 5427 (2007).
- Owens et al. [2015] A. G. Owens, S. Brühl, S. Simison, C. Forsich, and D. Heim, Comparison of tribological properties of stainless steel with hard and soft dlc coatings, Procedia Materials Science 9, 246 (2015).
- Luo et al. [2007] J. K. Luo, Y. Q. Fu, H. R. Le, J. A. Williams, S. M. Spearing, and W. I. Milne, Diamond and diamond-like carbon mems, Journal of Micromechanics and Microengineering 17, S147 (2007).
- Kasiorowski et al. [2020] T. Kasiorowski, J. Lin, P. Soares, C. Lepienski, C. Neitzke, G. de Souza, and R. Torres, Microstructural and tribological characterization of dlc coatings deposited by plasma enhanced techniques on steel substrates, Surface and Coatings Technology 389, 125615 (2020).
- Weiser et al. [1995] P. S. Weiser, S. Prawer, R. R. Manory, A. Hoffman, P. J. Evans, and P. J. Paterson, Chemical vapour deposition of diamond onto steel: the effect of a ti implant layer, Surface and Coatings Technology 71, 167 (1995).
- Chang and Wang [2001] C.-L. Chang and D.-Y. Wang, Microstructure and adhesion characteristics of diamond-like carbon films deposited on steel substrates, Diamond and Related Materials 10, 1528 (2001).
- Guo et al. [2010] H. Guo, Y. Qi, and X. Li, Adhesion at diamond/metal interfaces: A density functional theory study, Journal of Applied Physics 107, 033722 (2010).
- Wang et al. [2021] K. Wang, H. Zhou, K. Zhang, X. Liu, X. Feng, Y. Zhang, G. Chen, and Y. Zheng, Effects of ti interlayer on adhesion property of dlc films: A first principle study, Diamond and Related Materials 111, 108188 (2021).
- Hagarová et al. [2024] M. Hagarová, G. Baranová, M. Heželová, M. Truchlý, M. Vojtko, O. Petruš, and D. Csík, High-temperature mechanical and tribological performance of w-dlc coating with cr interlayer on x40crmov5-1 hot work tool steel, Coatings 14, 10.3390/coatings14080971 (2024).
- Ilic et al. [2019] E. Ilic, A. Pardo, T. Suter, S. Mischler, P. Schmutz, and R. Hauert, A methodology for characterizing the electrochemical stability of dlc coated interlayers and interfaces, Surface and Coatings Technology 375, 402 (2019).
- Bentzon et al. [1994] M. Bentzon, K. Mogensen, J. B. Hansen, C. Barholm-Hansen, C. Triholt, P. Holiday, and S. Eskildsen, Metallic interlayers between steel and diamond-like carbon, Surface and Coatings Technology 68-69, 651 (1994).
- Vetter [2014] J. Vetter, 60years of dlc coatings: Historical highlights and technical review of cathodic arc processes to synthesize various dlc types, and their evolution for industrial applications, Surface and Coatings Technology 257, 213 (2014).
- Grill [1999] A. Grill, Diamond-like carbon: state of the art, Diamond and Related Materials 8, 428 (1999).
- Gershman et al. [2022] I. Gershman, A. Mironov, A. Mezrin, E. Torskaya, T. Kuznetsova, V. Lapitskaya, and A. Rogachev, Effect of sp3–sp2 transformation on the wear rate of the dlc coating, Lubricants 10, 10.3390/lubricants10050085 (2022).
- Yamamoto et al. [2012] S. Yamamoto, A. Kawana, H. Ichimura, and C. Masuda, Relationship between tribological properties and sp3/sp2 structure of nitrogenated diamond-like carbon deposited by plasma cvd, Surface and Coatings Technology 210, 1 (2012).
- Li et al. [2020] C. Li, L. Huang, and J. Yuan, Effect of sp3 content on adhesion and tribological properties of non-hydrogenated dlc films, Materials 13, 10.3390/ma13081911 (2020).
- Cheng et al. [2024] C. Cheng, X. Sun, X. He, E. Xu, Q. Wang, R. Cao, W. Shen, F. Dong, Y. Guo, Z. Zhang, and S. Liu, Atomic and electronic properties of the metal/diamond (100) interfaces by first-principles calculations, Surfaces and Interfaces 52, 104916 (2024).
- Ichibha et al. [2018] T. Ichibha, K. Hongo, I. Motochi, N. Makau, G. Amolo, and R. Maezono, Adhesion of electrodes on diamond (111) surface: A dft study, Diamond and Related Materials 81, 168 (2018).
- Wang and Smith [2001] X.-G. Wang and J. R. Smith, Copper/diamond adhesion and hydrogen termination, Phys. Rev. Lett. 87, 186103 (2001).
- Zhang et al. [2022] S. Zhang, L. Rao, W. Shao, Q. He, X. Xing, Y. Zhou, and Q. Yang, Electronic, tensile, and sliding characteristics of the c/ti interface: A first-principles study, Langmuir 38, 15113 (2022).
- Zur and McGill [1984] A. Zur and T. C. McGill, Lattice match: An application to heteroepitaxy, Journal of Applied Physics 55, 378 (1984).
- Pierre et al. [2014] M. D. L. Pierre, M. Bruno, C. Manfredotti, F. Nestola, M. Prencipe, and C. Manfredotti, The (100), (111) and (110) surfaces of diamond: an ab initio b3lyp study, Molecular Physics 112, 1030 (2014).
- Stekolnikov et al. [2002] A. A. Stekolnikov, J. Furthmüller, and F. Bechstedt, Absolute surface energies of group-iv semiconductors: Dependence on orientation and reconstruction, Phys. Rev. B 65, 115318 (2002).
- Bechstedt [2003] F. Bechstedt, Principles of Surface Physics, 1st ed., Advanced Texts in Physics (Springer, Berlin, Heidelberg, 2003).
- Losi et al. [2023] G. Losi, O. Chehaimi, and M. C. Righi, Tribchem: A software for the first-principles, high-throughput study of solid interfaces and their tribological properties, Journal of Chemical Theory and Computation 19, 5231 (2023).
- Kresse and Hafner [1993] G. Kresse and J. Hafner, Ab initio molecular dynamics for liquid metals, Phys. Rev. B 47, 558 (1993).
- Kresse and Furthmüller [1996] G. Kresse and J. Furthmüller, Efficiency of ab-initio total energy calculations for metals and semiconductors using a plane-wave basis set, Comput. Mater. Sci. 6, 15 (1996).
- Kresse and Furthmüller [1996] G. Kresse and J. Furthmüller, Efficient iterative schemes for ab initio total-energy calculations using a plane-wave basis set, Phys. Rev. B 54, 11169 (1996).
- Jain et al. [2015] A. Jain, S. P. Ong, W. Chen, B. Medasani, X. Qu, M. Kocher, M. Brafman, G. Petretto, G.-M. Rignanese, G. Hautier, D. Gunter, and K. A. Persson, Fireworks: a dynamic workflow system designed for high-throughput applications, Concurrency and Computation: Practice and Experience 27, 5037 (2015).
- Mathew et al. [2017] K. Mathew, J. H. Montoya, A. Faghaninia, S. Dwarakanath, M. Aykol, H. Tang, I. heng Chu, T. Smidt, B. Bocklund, M. Horton, J. Dagdelen, B. Wood, Z.-K. Liu, J. Neaton, S. P. Ong, K. Persson, and A. Jain, Atomate: A high-level interface to generate, execute, and analyze computational materials science workflows, Computational Materials Science 139, 140 (2017).
- [42] Custodian, https://github.com/materialsproject/custodian.
- Ong et al. [2013] S. P. Ong, W. D. Richards, A. Jain, G. Hautier, M. Kocher, S. Cholia, D. Gunter, V. L. Chevrier, K. A. Persson, and G. Ceder, Python materials genomics (pymatgen): A robust, open-source python library for materials analysis, Computational Materials Science 68, 314 (2013).
- Mathew et al. [2016] K. Mathew, A. K. Singh, J. J. Gabriel, K. Choudhary, S. B. Sinnott, A. V. Davydov, F. Tavazza, and R. G. Hennig, Mpinterfaces: A materials project based python tool for high-throughput computational screening of interfacial systems, Computational Materials Science 122, 183 (2016).
- Wolloch et al. [2019] M. Wolloch, G. Losi, M. Ferrario, and M. C. Righi, High-throughput screening of the static friction and ideal cleavage strength of solid interfaces, Scientific Reports 9, https://doi.org/10.1038/s41598-019-49907-2 (2019).
- Restuccia et al. [2023] P. Restuccia, G. Losi, O. Chehaimi, M. Marsili, and M. C. Righi, High-throughput first-principles prediction of interfacial adhesion energies in metal-on-metal contacts, ACS Applied Materials & Interfaces 15, 19624 (2023).
- Perdew et al. [1996] J. P. Perdew et al., Generalized gradient approximation made simple, Phys. Rev. Lett. 77, 3865 (1996).
- Jain et al. [2013] A. Jain, S. P. Ong, G. Hautier, W. Chen, W. D. Richards, S. Dacek, S. Cholia, D. Gunter, D. Skinner, G. Ceder, and K. A. Persson, Commentary: The materials project: A materials genome approach to accelerating materials innovation, APL Materials 1, 011002 (2013).
- Birch [1947] F. Birch, Finite elastic strain of cubic crystals, Phys. Rev. 71, 809 (1947).
- Damiani et al. [2024] E. Damiani, M. Marsili, and M. C. Righi, Tuning the adhesion of diamond/copper interfaces through surface chemical modifications and reconstruction, Carbon 230, 119555 (2024).
- Israelachvili [2011] J. N. Israelachvili, in Intermolecular and Surface Forces, edited by J. N. Israelachvili (Academic Press, San Diego, 2011) 3rd ed., pp. 191–204.
- Young and Goodwin [2011] C. A. Young and A. L. Goodwin, Applications of pair distribution function methods to contemporary problems in materials chemistry, J. Mater. Chem. 21, 6464 (2011).
- Billinge [2008] S. J. Billinge, Nanoscale structural order from the atomic pair distribution function (pdf): There’s plenty of room in the middle, Journal of Solid State Chemistry 181, 1695 (2008).
- Fernandez et al. [2013] M. Fernandez, N. R. Trefiak, and T. K. Woo, Atomic property weighted radial distribution functions descriptors of metal–organic frameworks for the prediction of gas uptake capacity, The Journal of Physical Chemistry C 117, 14095 (2013).
- Kajita and Righi [2016] S. Kajita and M. Righi, A fundamental mechanism for carbon-film lubricity identified by means of ab initio molecular dynamics, Carbon 103, 193 (2016).
- Pacini et al. [2024] A. Pacini, M. Ferrario, S. Loehle, and M. Righi, dvancing tribological simulations of carbon-based lubricants with active learning and machine learning molecular dynamics, Eur. Phys. J. Plus 139, https://doi.org/10.1140/epjp/s13360-024-05348-z (2024).
- Marsili et al. [2005] M. Marsili, O. Pulci, F. Bechstedt, and R. Del Sole, Electronic structure of the surface: Solution by self-consistent many-body calculations, Phys. Rev. B 72, 115415 (2005).
- Kabedev et al. [2024] A. Kabedev, S. Hossain, and P. Larsson, Molecular dynamics simulations as a tool to understand drug solubilization in pharmaceutical systems, in Comprehensive Computational Chemistry (First Edition), edited by M. Yáñez and R. J. Boyd (Elsevier, Oxford, 2024) first edition ed., pp. 865–885.
- Wolloch et al. [2018] M. Wolloch et al., Interfacial charge density and its connection to adhesion and frictional forces, Phys. Rev. Lett. 121 (2018).